# Magnetic properties of Cr8 and V8 molecular rings from ab initio calculations

### v1

@@ V Discussion
The agreement with experimental results on Cr8’s $J^{\left(1\right)}$ is further enhanced by using the extended DFT+U+V scheme with on-site and inter-site interactions (between TM magnetic centers and neighbor anions). While this result probably confirms the role of the intermediate anions orbitals in the superexchange mechanism, the overall qualitative picture is the same as with DFT+U.
@@
This behavior of V8 is rather unusual. Our preliminary calculations, where V centers are substituted with other TMs, consistently stabilize an AFM ground state. The V8 peculiarity seems to question the validity of the superexchange theory (also challenged by the presence of multiple anionic links between neighbor TM centers) and of the Goodenough-Kanamori-Anderson rules. However, it is important to note that in their simpler form, superexchange mechanisms are often referred to half-filling conditions. These mechanisms are derived from second-order perturbation theory applied to an Hubbard model with one orbital and one electron per site (half-filling). This condition is not satisfied by V8 that, in its 3+ oxidation state, hosts the two outermost electrons in a group of three quasi-degenerate $d$ states (this would be the $t_{2g}$ multiplet if the ligands’ octahedron around it had a perfect cubic symmetry). We believe that this less-than-half-filling condition of V3+ is what makes V8 to be FM. It is quite well established in literature that at less-than-half filling, a strong enough on-site repulsion between electrons ($U$ or $U-J$ for parallel spins) or, equivalently, a weak hopping amplitude ($t$), can cause a FM coupling between neighboring spins, mediated by the intermediate anion [102, 103].
To directly verify this point, we devised a simple Hubbard model with less-than-half-filled orbitals on the metallic centers, which we solved exactly to study the ground state energy as a function of $J/t$ ($J$ being the Hund’s exchange coupling for electrons on the same atomic site).
Figure 7: A schematic representation of the simple model adopted to study the superexchange coupling between less-than-half occupied metallic sites (M), as mediated by anions between them. $t$ is the hopping amplitude between metallic and anionic states; $\epsilon$ is the site energy of anionic states; $U$ and $U-J$ represent the on-site repulsion between electrons on the same metal with antiparallel or parallel spin, respectively.
As sketched in Fig. 7, the model contains two lateral transition-metal sites (indicated with M, each contributing two states) and two anionic sites (A) between them, each contributing one state. The total number of electrons is six, with anions levels guaranteed to be doubly-occupied by a negative site energy $\epsilon$, so that only one electron is present on each metallic site. This realizes a quarter-filling condition. The repulsion between electrons on the same metallic site is the Hubbard $U$possibly corrected by the exchange $J$ ($U$ or $U-J$ for antiparallel and parallel spins, respectively). The hybridization between metallic and anionic states is promoted by the hopping parameter $t$. The resulting 6-states Hubbard Hamiltonian can then be diagonalized exactly, e.g., using the HPHI code [104, 105].
Figure 8: Energy difference between the triplet and singlet ground states of the less-than-half-filled Hubbard model described in the text. All quantities are in units of the hopping amplitude $t$.
@@
This observation confirms the idea that, for strong enough on-site exchange interactions, less-than-half-filled metallic states can couple ferromagnetically, even for geometries that would otherwise promote AFM configurations. Therefore, the anomalous ground state of V8 can, at least in part, be traced back to the lower filling of its $d$ states.
@@ VI Conclusions
The anomalous behavior of V8 can be rationalized based on the less-than-half filling condition of the $d$ states submanifold hosting the valence electrons responsible for the localized moments, as qualitatively demonstrated by the exact diagonalization of a simple Hubbard model mimicking this situation.
@@ VII Acknowledgements
B. M. M. and M. C. acknowledge support from NQSTI within the PNRR MUR project PE0000023-NQSTI. Part of the calculations of this work were realized thanks to the high-performance computing resources and support made available by CINECA (through awards within the ISCRA initiative). Moreover, via A.F. membership of the UK’s HEC Materials Chemistry Consortium, funded by EPSRC (EP/X035859), this work also used the ARCHER2 UK National Supercomputing Service (http://www.archer2.ac.uk).

### v2

@@ V Discussion
The agreement with experimental results on Cr8’s $J^{\left(1\right)}$ is further enhanced by using the extended DFT+U+V scheme with on-site and inter-site interactions between TM magnetic centers and neighbor anions. While this result probably confirms the role of the intermediate anions orbitals in the superexchange mechanism, the overall qualitative picture is the same as with DFT+U.
@@
This behavior of V8 is rather unusual. Our preliminary calculations, where V centers are substituted with other TMs, consistently stabilize an AFM ground state. The V8 peculiarity seems to question the validity of the superexchange theory (also challenged by the presence of multiple anionic links between neighbor TM centers) and of the Goodenough-Kanamori-Anderson rules. However, it is important to note that in their simpler form, superexchange mechanisms are often referred to half-filling conditions. These mechanisms are derived from second-order perturbation theory applied to an Hubbard model with one orbital and one electron per site (half-filling). This condition is not satisfied by V8 that, in its 3+ oxidation state, hosts the two outermost electrons in a group of three quasi-degenerate $d$ states, namely the $t_{2g}$ multiplet if the ligands’ octahedron around it had a perfect cubic symmetry. We believe that this less-than-half-filling condition of V3+ is at the origin of V8’s FM ground state. In fact, it is quite well established in literature that at less-than-half filling, a strong enough on-site repulsion between electrons ($U$ or $U-J$ for parallel spins) or, equivalently, a weak hopping amplitude ($t$), can cause a FM coupling between neighboring spins, mediated by the intermediate anion [105, 106].
Figure 7: A modified V8 ring highlighting the presence of one extra H bound to each F atom in the internal space of the molecule.
In order to verify this idea, we have performed a series of tests on modified V8 or Cr8 rings that feature a different anionic group at the F site: either an H-coordinated F or an O, so as to change the number of $d$ electrons on the magnetic centers. Fig. 7 shows, as an example, the ring with an extra H bound to each internal F. It is important to note that besides the modification of the anionic center, no extra geometrical optimization has been performed. The evaluation of the effective $J$’s was achieved by comparing the FM and AFM configurations energies, computed with LDA+U.
System Anion Av. mag. ($\mu_{B}$) $J$ (meV) stable
Cr8 F 3 0.843 AFM
O 2 -148.188 FM
V8 F 2 -0.917 FM
F-H 3 7.576 AFM
O 1 -22.456 FM
@@
Table 5 compares the average absolute magnetization per TM center and the $J$’s for Cr8 with F replaced by O, and for V8 with F replaced by F-H and O. “Standard” values, obtained with F on the internal anionic site are also reported to ease the comparison (these are the same as in Table 3). In all cases the substitution of F anions with F-H groups (with single O) leads to a reduction (oxidation) of the TM centers that couple to a consistent unitary increase (decrease) in the absolute magnetization of the TMs. The V centers reduction brings 3 electrons to the $3d$ states (like Cr in Cr8) which achieves a 2+ state and a half filling condition for the occupied submanifold. Correspondingly, the $J$ becomes positive and large (7.576 meV) which stabilizes an AFM configuration. On the contrary, with O replacing F, the TM centers get a 4+ state that promotes a further departure from half-filling. With less than half-filled $3d$ submanifolds both systems exhibit negative $J$’s that make FM spin configurations more stable. In this case the $J$’s are very large in absolute value (-22.456 meV and -148.188 meV for V8 and Cr8, respectively). Due to the high energy required to reach a 4+ oxidation state the hole introduced in the structure is actually spread over the rings with a consequent severe overestimation of the computed couplings (similarly to what is obtained with LDA or GGA). However, the main focus of this discussion is the sign of the magnetic couplings. Our conjecture is thus overall confirmed: magnetic centers with less-than-half-filled sub-groups of $3d$ states can couple ferromagnetically, even for geometries that would otherwise promote AFM configurations. Therefore, the anomalous ground state of V8 can, at least in part, be traced back to the lower filling of its $d$ states.
@@ VI Conclusions
The anomalous behavior of V8 can be traced back to the less-than-half filling condition of the $d$ states submanifold of V3+ hosting the valence electrons responsible for the ion’s localized moments. This idea has been proved by evaluating the effective exchange coupling in modified systems where the number of $d$ electrons on the TM ions was artificially changed by either coordinating F ions with extra H atoms or substituting them with O anions.
@@ VII Acknowledgements
The authors are thankful to Prof. L. Kronik for a preliminary discussion on the Cr8 system and for providing its initial structure. B. M. M. and M. C. acknowledge support from NQSTI within the PNRR MUR project PE0000023-NQSTI. Part of the calculations of this work were realized thanks to the high-performance computing resources and support made available by CINECA (through awards within the ISCRA initiative). Moreover, via A.F. membership of the UK’s HEC Materials Chemistry Consortium, funded by EPSRC (EP/X035859), this work also used the ARCHER2 UK National Supercomputing Service (http://www.archer2.ac.uk).